# Supersonic–sonic patch solution for the two-dimensional Euler equations with a van der Waals equation of state 11footnotemark: 1

### v1

@@ 2.4 Sonic-supersonic boundary data and the main conclusion in terms of $(\theta,\bar{\omega})$
$\begin{split}&\psi^{\prime}(\xi),{\hat{\bar{\omega}}}(\xi)\in C^{2}([\xi_{1},\xi_{2}]),\\ &\psi^{\prime}(\xi_{2})<0,\psi^{\prime\prime}(\xi_{2})<0,\hat{\bar{\omega}}^{\prime}(\xi_{2})>0.\end{split}$
@@ 3.2 The degenerate system and uniform bounds
$\begin{cases}\partial_{-}\overline{X}=\frac{A(\kappa(\tau))\overline{X}}{\sqrt{1-t^{2}}\mathcal{T}_{2}}\left\{\frac{1}{A(\kappa(\tau))\mu^{2}\sqrt{1+\kappa(\tau)(1-t^{2})}}\cdot\frac{\overline{X}-\overline{Y}}{2t}+\frac{\sqrt{1-t^{2}}(3+4\kappa(1-t^{2}))}{c(1+\kappa(1-t^{2}))}t^{2}\right\},\\ \partial_{+}\overline{Y}=-\frac{A(\kappa(\tau))\overline{Y}}{\sqrt{1-t^{2}}\mathcal{T}_{1}}\left\{\frac{1}{A(\kappa(\tau))\mu^{2}\sqrt{1+\kappa(\tau)(1-t^{2})}}\cdot\frac{\overline{X}-\overline{Y}}{2t}+\frac{\sqrt{1-t^{2}}(3+4\kappa(1-t^{2}))}{c(1+\kappa(1-t^{2}))}t^{2}\right\},\end{cases}$ (3.8)
@@
$\begin{split}f(t,\tau)&=\frac{1}{(1+\kappa(\tau)-\kappa(\tau)t^{2})(1-t^{2})^{3/2}}\;,\;g(t,\tau)=-\frac{(1-t^{2})^{2}f}{c},\\ \mathcal{T}_{1}&=\overline{Y}\sqrt{1+\kappa(\tau)(1-t^{2})}-\frac{t(1-t^{2})A(\kappa(\tau))}{c},\;\\ \mathcal{T}_{2}&=\overline{X}\sqrt{1+\kappa(\tau)(1-t^{2})}+\frac{t(1-t^{2})A(\kappa(\tau))}{c}.\end{split}$ (3.9)
@@ 3.4 Existence of global solution in $\Omega$
$\begin{split}H_{11}&=-\frac{\sqrt{1-t^{2}}\widetilde{Y}}{2c(1+\kappa-\kappa t^{2})(1+tg\widetilde{Y})},\;\;H_{21}=-\frac{\sqrt{1-t^{2}}\widetilde{X}}{2c(1+\kappa-\kappa t^{2})(1-tg\widetilde{X})},\\ H_{12}&=\frac{(\widetilde{X}-\widetilde{Y})(1+2\kappa-\kappa t^{2})}{2(1-t^{2})\sqrt{1+\kappa-\kappa t^{2}}T_{3}}-\frac{f\sqrt{1-t^{2}}}{1+tg\widetilde{Y}}\left(\frac{2t^{2}\sqrt{1-t^{2}}\widetilde{X}\widetilde{Y}}{c}-\left(\tau\kappa^{\prime}(\tau)-1-2\kappa(\tau)+2t^{2}\kappa(\tau)\right)\widetilde{X}t\right),\\ H_{22}&=-\frac{(\widetilde{X}-\widetilde{Y})(1+2\kappa-\kappa t^{2})}{2(1-t^{2})\sqrt{1+\kappa-\kappa t^{2}}T_{4}}+\frac{f\sqrt{1-t^{2}}}{1-tg\widetilde{X}}\left(\frac{2t^{2}\sqrt{1-t^{2}}\widetilde{X}\widetilde{Y}}{c}-\left(\tau\kappa^{\prime}(\tau)-1-2\kappa(\tau)+2t^{2}\kappa(\tau)\right)\widetilde{Y}t\right).\end{split}$ (3.24)
@@
$\begin{split}G_{11}&=-\frac{1}{2t}+H_{11}+tH_{12\widetilde{X}},\;\;G_{21}=\frac{1}{2t}-H_{21}+tH_{22\widetilde{Y}},\\ G_{12}&=\left(-\frac{1}{2t}-H_{11}+(\widetilde{X}-\widetilde{Y})H_{11\widetilde{Y}}+tH_{12\widetilde{Y}}\right)\cdot\left(-\frac{\widetilde{X}-\widetilde{Y}}{2t}+H_{21}(\widetilde{X}-\widetilde{Y})+H_{22}t\right),\\ G_{22}&=\left(-\frac{1}{2t}+H_{21}+(\widetilde{X}-\widetilde{Y})H_{21\widetilde{X}}+tH_{22\widetilde{X}}\right)\cdot\left(\frac{\widetilde{X}-\widetilde{Y}}{2t}+H_{11}(\widetilde{X}-\widetilde{Y})+H_{12}t\right).\end{split}$
@@ 3.5 Regularity of the solution
$\begin{split}\begin{cases}\tilde{\partial}_{+}\Xi&=\frac{\tilde{\partial}_{-}\tilde{\Lambda}_{+}-\tilde{\partial}_{+}\tilde{\Lambda}_{-}}{\tilde{\Lambda}_{+}-\tilde{\Lambda}_{-}}\Xi+(\tilde{\Lambda}_{+}-\tilde{\Lambda}_{-})(\tilde{\partial}_{+}\widetilde{X})_{z},\\ \tilde{\partial}_{-}\Theta&=\frac{\tilde{\partial}_{-}\tilde{\Lambda}_{+}-\tilde{\partial}_{+}\tilde{\Lambda}_{-}}{\tilde{\Lambda}_{+}-\tilde{\Lambda}_{-}}\Theta+(\tilde{\Lambda}_{+}-\tilde{\Lambda}_{-})(\tilde{\partial}_{-}\widetilde{Y})_{z}.\end{cases}\end{split}$ (3.34)
@@
$\begin{split}f_{1}(z,t)=&\frac{1}{2t}-\frac{(\kappa(\tau)+1)g\widetilde{Y}-t(1+2\kappa(\tau)-t^{2}\kappa(\tau))(1+tg\widetilde{Y})}{2(1-t^{2})(1+tg\widetilde{Y})(1+\kappa(\tau)-t^{2}\kappa(\tau))}\\ &+\frac{f\sqrt{1-t^{2}}}{1+tg\widetilde{Y}}\left(t(\tau\kappa^{\prime}(\tau)-1-2\kappa(\tau)+2t^{2}\kappa(\tau))-\frac{2t^{2}f\sqrt{1-t^{2}}\widetilde{Y}}{c}\right),\\ f_{2}(z,t)=&-\frac{1}{2t}+\frac{(\kappa(\tau)+1)g\widetilde{Y}-t(1+2\kappa(\tau)-t^{2}\kappa(\tau))(1+tg\widetilde{Y})}{2(1-t^{2})(1+tg\widetilde{Y})(1+\kappa(\tau)-t^{2}\kappa(\tau))}-\frac{2t^{2}f(1-t^{2})\widetilde{X}}{c(1+tg\widetilde{Y})}\\ &-\frac{fg\sqrt{1-t^{2}}}{(1+tg\widetilde{Y})^{2}}\left(\frac{\widetilde{X}-\widetilde{Y}}{2\mu^{2}}+(\tau\kappa^{\prime}(\tau)-1-2\kappa(\tau)+2t^{2}\kappa(\tau))t^{2}\widetilde{X}-\frac{2t^{3}f\sqrt{1-t^{2}}\widetilde{X}\widetilde{Y}}{c}\right),\\ f_{3}(z,t)=&(\widetilde{Y}-\widetilde{X})\left(\frac{\kappa^{\prime}(\tau)\tau_{z}}{1+\kappa(\tau)}+\frac{f_{z}}{f(1+tg\widetilde{Y})}\right)\left(\frac{(\kappa(\tau)+1)g\widetilde{Y}-t(1+2\kappa(\tau)-t^{2}\kappa(\tau))(1+tg\widetilde{Y})}{2(1+\kappa(\tau)-t^{2}\kappa(\tau))(1-t^{2})(1+tg\widetilde{Y})}\right)\\ &+\frac{\widetilde{X}-\widetilde{Y}}{2t}\left(\frac{\kappa^{\prime}(\tau)\tau_{z}}{1+\kappa(\tau)}+\frac{f_{z}}{f(1+tg\widetilde{Y})}\right)+\frac{f\sqrt{1-t^{2}}}{1+tg\widetilde{Y}}\left(-\frac{2t^{2}\sqrt{1-t^{2}}\widetilde{X}\widetilde{Y}}{\tau\kappa(\tau)c}+\widetilde{X}t[\kappa^{\prime}(\tau)(2t^{2}-1)+\tau\kappa^{\prime\prime}(\tau)]\right)\tau_{z}\\ &+\frac{f_{z}\sqrt{1-t^{2}}}{(1+tg\widetilde{Y})^{2}}\left(-\frac{2t^{2}\sqrt{1-t^{2}}\widetilde{X}\widetilde{Y}}{c}+\widetilde{X}t[\tau\kappa^{\prime}(\tau)-1-2\kappa(\tau)+t^{2}2\kappa(\tau))]\right)\\ &+\frac{fg\widetilde{Y}\sqrt{1-t^{2}}}{c(1+tg\widetilde{Y})^{2}}\left(\frac{\widetilde{X}-\widetilde{Y}}{2\mu^{2}}-\frac{2t^{3}\sqrt{1-t^{2}}\widetilde{X}\widetilde{Y}}{c}+\widetilde{X}t^{2}[\tau\kappa^{\prime}(\tau)-1-2\kappa(\tau)+t^{2}\kappa(\tau))]\right)c_{z},\\ g_{1}(z,t)=&\frac{1}{2t}+\frac{(\kappa(\tau)+1)g\widetilde{X}+t(1+2\kappa(\tau)-t^{2}\kappa(\tau))(1-tg\widetilde{X})}{2(1-t^{2})(1-tg\widetilde{X})(1+\kappa(\tau)-t^{2}\kappa(\tau))}\\ &+\frac{f\sqrt{1-t^{2}}}{1-tg\widetilde{X}}\left(t(\tau\kappa^{\prime}(\tau)-1-2\kappa(\tau)+2t^{2}\kappa(\tau))+\frac{2t^{2}f\sqrt{1-t^{2}}\widetilde{X}}{c}\right),\\ g_{2}(z,t)=&-\frac{1}{2t}-\frac{(\kappa(\tau)+1)g\widetilde{X}+t(1+4\kappa(\tau)-2t^{2}\kappa(\tau))(1-tg\widetilde{X})}{2(1-t^{2})(1-tg\widetilde{X})(1+\kappa(\tau)-t^{2}\kappa(\tau))}-\frac{2t^{2}f(1-t^{2})\widetilde{Y}}{c(1-tg\widetilde{X})}\\ &-\frac{fg\sqrt{1-t^{2}}}{(1-tg\widetilde{X})^{2}}\left(\frac{\widetilde{X}-\widetilde{Y}}{2\mu^{2}}-(\tau\kappa^{\prime}(\tau)-1-2\kappa(\tau)+2t^{2}\kappa(\tau))t^{2}\widetilde{Y}-\frac{2t^{3}f\sqrt{1-t^{2}}\widetilde{X}\widetilde{Y}}{c}\right),\\ g_{3}(z,t)=&(\widetilde{Y}-\widetilde{X})\left(\frac{\kappa^{\prime}(\tau)\tau_{z}}{1+\kappa(\tau)}+\frac{f_{z}}{f(1-tg\widetilde{X})}\right)\left(\frac{(\kappa(\tau)+1)g\widetilde{X}+t(1+4\kappa(\tau)-2t^{2}\kappa(\tau))(1-tg\widetilde{X})}{2(1+\kappa(\tau)-t^{2}\kappa(\tau))(1-t^{2})(1-tg\widetilde{X})}\right)\\ &+\frac{\widetilde{Y}-\widetilde{X}}{2t}\left(\frac{\kappa^{\prime}(\tau)\tau_{z}}{1+\kappa(\tau)}+\frac{f_{z}}{f(1-tg\widetilde{X})}\right)+\frac{f\sqrt{1-t^{2}}}{1-tg\widetilde{X}}\left(\frac{2t^{2}\sqrt{1-t^{2}}\widetilde{X}\widetilde{Y}}{\tau\kappa(\tau)c}+\widetilde{Y}t[\kappa^{\prime}(\tau)(2t^{2}-1)+\tau\kappa^{\prime\prime}(\tau)]\right)\tau_{z}\\ &+\frac{f_{z}\sqrt{1-t^{2}}}{(1-tg\widetilde{X})^{2}}\left(\frac{2t^{2}\sqrt{1-t^{2}}\widetilde{X}\widetilde{Y}}{c}+\widetilde{Y}t[\tau\kappa^{\prime}(\tau)-1-2\kappa(\tau)+t^{2}2\kappa(\tau))]\right)\\ &+\frac{fg\widetilde{X}\sqrt{1-t^{2}}}{c(1-tg\widetilde{X})^{2}}\left(\frac{\widetilde{X}-\widetilde{Y}}{2\mu^{2}}-\frac{2t^{3}\sqrt{1-t^{2}}\widetilde{X}\widetilde{Y}}{c}+\widetilde{Y}t^{2}[\tau\kappa^{\prime}(\tau)-1-2\kappa(\tau)+t^{2}\kappa(\tau))]\right)c_{z}.\end{split}$
@@
$\begin{split}\tilde{\partial}_{+}\Xi=\left(\frac{5}{2t}+\hat{f}_{1}(z,t)\right)\Xi+\hat{f}_{2}(z,t)\frac{\Theta}{2t}+\hat{f}_{3}(z,t)\frac{t}{2},\\ \tilde{\partial}_{-}\Theta=\left(\frac{5}{2t}+\hat{g}_{1}(z,t)\right)\Xi+\hat{g}_{2}(z,t)\frac{\Xi}{2t}+\hat{g}_{3}(z,t)\frac{t}{2},\end{split}$ (3.36)
@@
$\begin{split}\hat{f}_{1}&=h+\left(f_{1}-\frac{1}{2t}\right),\;\hat{f}_{2}=2tf_{2},\;\hat{f}_{3}=\frac{2}{t}(\tilde{\Lambda}_{+}-\tilde{\Lambda}_{-})f_{3},\\ \hat{g}_{1}&=h+\left(g_{1}-\frac{1}{2t}\right),\;\hat{g}_{2}=2tg_{2},\;\hat{g}_{3}=\frac{2}{t}(\tilde{\Lambda}_{+}-\tilde{\Lambda}_{-})g_{3}.\end{split}$
@@
$\begin{split}\left(\frac{\hat{c}\hat{f}t}{\hat{b}+t\hat{g}}-\frac{\hat{c}\sqrt{1+(\psi^{\prime})^{2}}}{\hat{\bar{\omega}}^{2}\hat{\bar{\omega}}^{\prime}}\right)\cdot\tilde{X}_{z}|_{\widehat{M^{\prime}P^{\prime}}}=&\frac{\hat{b}-\hat{a}}{2t^{2}\hat{a}\hat{b}}+\frac{\hat{b}-\hat{a}}{\hat{a}\hat{b}t}\hat{H}_{11}+\hat{H}_{12}+\frac{\hat{d}}{t\hat{a}^{2}}-\frac{t\hat{d}^{\prime}}{\hat{\bar{\omega}}\hat{\bar{\omega}}^{\prime}\hat{a}^{2}}-\frac{1}{\hat{\bar{\omega}}\hat{\bar{\omega}}^{\prime}\hat{a}^{2}}\left(\frac{\cos\hat{\theta}\psi^{\prime\prime}}{\kappa(\hat{\tau})\bar{\omega}(1+(\psi)^{\prime 2})}\right)^{\prime}\\ =&-\frac{2\hat{d}^{2}}{\hat{a}^{2}\hat{b}}-\frac{2\hat{d}}{\hat{a}\hat{b}}\hat{H}_{11}+\hat{H}_{12}-\frac{t\hat{d}^{\prime}}{\hat{\bar{\omega}}\hat{\bar{\omega}}^{\prime}\hat{a}^{2}}-\frac{1}{\hat{\bar{\omega}}\hat{\bar{\omega}}^{\prime}\hat{a}^{2}}\left(\frac{\cos\hat{\theta}\psi^{\prime\prime}}{\kappa(\hat{\tau})\bar{\omega}(1+(\psi)^{\prime 2})}\right)^{\prime},\end{split}$
@@
$\begin{split}\tilde{\partial}_{+}\left(t^{-\frac{1}{2}-\nu}\widetilde{\Xi}\right)&=t^{-\frac{3}{2}-\nu}\left(\hat{f}_{1}t\widetilde{\Xi}+\frac{\widetilde{\Theta}}{2}\hat{f}_{2}+\frac{t^{\nu}}{2}\hat{f}_{3}\right),\\ \tilde{\partial}_{-}\left(t^{-\frac{1}{2}-\nu}\widetilde{\Theta}\right)&=t^{-\frac{3}{2}-\nu}\left(\hat{g}_{1}t\widetilde{\Theta}+\frac{\widetilde{\Xi}}{2}\hat{g}_{2}+\frac{t^{\nu-1}}{2}\hat{g}_{3}\right).\end{split}$ (3.39)

### v2

@@ 2.4 Sonic-supersonic boundary data and the main conclusion in terms of $(\theta,\bar{\omega})$
$\begin{split}&\psi^{\prime}(\xi),{\hat{\bar{\omega}}}(\xi)\in C^{2}([\xi_{1},\xi_{2}]),\\ &\psi^{\prime}(\xi_{2})<0,\psi^{\prime\prime}(\xi_{2})<0,\hat{\bar{\omega}}^{\prime}(\xi_{2})>0.\end{split}$ (2.25)
@@ 3.2 The degenerate system and uniform bounds
$\begin{cases}\partial_{-}\overline{X}=-\frac{A(\kappa(\tau))\overline{X}}{\sqrt{1-t^{2}}\mathcal{T}_{1}}\left\{\frac{1}{A(\kappa(\tau))\mu^{2}\sqrt{1+\kappa(\tau)(1-t^{2})}}\cdot\frac{\overline{X}-\overline{Y}}{2t}+\frac{\sqrt{1-t^{2}}(3+4\kappa(1-t^{2}))}{c(1+\kappa(1-t^{2}))}t^{2}\right\},\\ \partial_{+}\overline{Y}=-\frac{A(\kappa(\tau))\overline{Y}}{\sqrt{1-t^{2}}\mathcal{T}_{2}}\left\{\frac{1}{A(\kappa(\tau))\mu^{2}\sqrt{1+\kappa(\tau)(1-t^{2})}}\cdot\frac{\overline{X}-\overline{Y}}{2t}+\frac{\sqrt{1-t^{2}}(3+4\kappa(1-t^{2}))}{c(1+\kappa(1-t^{2}))}t^{2}\right\},\end{cases}$ (3.8)
@@
$\begin{split}f(t,\tau)&=\frac{1}{(1+\kappa(\tau)-\kappa(\tau)t^{2})(1-t^{2})^{3/2}}\;,\;g(t,\tau)=-\frac{(1-t^{2})^{2}f}{c},\\ \mathcal{T}_{1}&=\frac{A(\kappa(\tau)(Y-tg)}{f(1-t^{2})},\;\mathcal{T}_{2}=\frac{A(\kappa(\tau))(X-tg)}{f(1-t^{2})}.\end{split}$ (3.9)
@@ 3.4 Existence of global solution in $\Omega$
[…]
@@
$\begin{split}G_{11}=&-\frac{1}{2t}+H_{11}+tH_{12\widetilde{X}},\;\;G_{21}=\frac{1}{2t}-H_{21}+tH_{22\widetilde{Y}},\\ G_{12}=&\left(-\frac{1}{2t}-H_{11}+(\widetilde{X}-\widetilde{Y})H_{11\widetilde{Y}}+tH_{12\widetilde{Y}}\right)\cdot\left(-\frac{\widetilde{X}-\widetilde{Y}}{2t}+H_{21}(\widetilde{X}-\widetilde{Y})+H_{22}t\right)-\frac{\widetilde{X}-\widetilde{Y}}{2t^{2}}(1-2t^{2}H_{11t})\\ &+\tilde{\Lambda}_{-}(tH_{12z}+(\widetilde{X}-\widetilde{Y})H_{11z})+H_{12}+tH_{12t},\\ G_{22}=&\left(-\frac{1}{2t}+H_{21}+(\widetilde{X}-\widetilde{Y})H_{21\widetilde{X}}+tH_{22\widetilde{X}}\right)\cdot\left(\frac{\widetilde{X}-\widetilde{Y}}{2t}+H_{11}(\widetilde{X}-\widetilde{Y})+H_{12}t\right)+\frac{\widetilde{X}-\widetilde{Y}}{2t^{2}}(1-2t^{2}H_{21t})\\ &+\tilde{\Lambda}_{+}(tH_{22z}+(\widetilde{X}-\widetilde{Y})H_{21z})+H_{22}+tH_{22t}.\end{split}$
@@ 3.5 Regularity of the solution
$\begin{split}\tilde{\partial}_{+}\Xi&=\frac{\tilde{\partial}_{-}\tilde{\Lambda}_{+}-\tilde{\partial}_{+}\tilde{\Lambda}_{-}}{\tilde{\Lambda}_{+}-\tilde{\Lambda}_{-}}\Xi+(\tilde{\Lambda}_{+}-\tilde{\Lambda}_{-})(\tilde{\partial}_{+}\widetilde{X})_{z},\\ \tilde{\partial}_{-}\Theta&=\frac{\tilde{\partial}_{-}\tilde{\Lambda}_{+}-\tilde{\partial}_{+}\tilde{\Lambda}_{-}}{\tilde{\Lambda}_{+}-\tilde{\Lambda}_{-}}\Theta+(\tilde{\Lambda}_{+}-\tilde{\Lambda}_{-})(\tilde{\partial}_{-}\widetilde{Y})_{z}.\end{split}$ (3.34)
@@
$\begin{split}f_{1}(z,t)=&\frac{1}{2t}-\frac{(\kappa(\tau)+1)g\widetilde{Y}-t(1+2\kappa(\tau)-t^{2}\kappa(\tau))(1+tg\widetilde{Y})}{2(1-t^{2})(1+tg\widetilde{Y})(1+\kappa(\tau)-t^{2}\kappa(\tau))}\\ &+\frac{f\sqrt{1-t^{2}}}{1+tg\widetilde{Y}}\left(t(\tau\kappa^{\prime}(\tau)-1-2\kappa(\tau)+2t^{2}\kappa(\tau))-\frac{2t^{2}\sqrt{1-t^{2}}\widetilde{Y}}{c}\right),\\ f_{2}(z,t)=&-\frac{1}{2t}+\frac{(\kappa(\tau)+1)g\widetilde{Y}-t(1+2\kappa(\tau)-t^{2}\kappa(\tau))(1+tg\widetilde{Y})}{2(1-t^{2})(1+tg\widetilde{Y})(1+\kappa(\tau)-t^{2}\kappa(\tau))}-\frac{2t^{2}f(1-t^{2})\widetilde{X}}{c(1+tg\widetilde{Y})}\\ &-\frac{fg\sqrt{1-t^{2}}}{(1+tg\widetilde{Y})^{2}}\left(\frac{\widetilde{X}-\widetilde{Y}}{2\mu^{2}(\tau)}+(\tau\kappa^{\prime}(\tau)-1-2\kappa(\tau)+2t^{2}\kappa(\tau))t^{2}\widetilde{X}-\frac{2t^{3}\sqrt{1-t^{2}}\widetilde{X}\widetilde{Y}}{c}\right),\\ f_{3}(z,t)=&(\widetilde{Y}-\widetilde{X})\left(\mu^{2}(\tau)\kappa^{\prime}(\tau)\tau_{z}+\frac{f_{z}}{f(1+tg\widetilde{Y})}\right)\left(\frac{(\kappa(\tau)+1)g\widetilde{Y}-t(1+2\kappa(\tau)-t^{2}\kappa(\tau))(1+tg\widetilde{Y})}{2(1+\kappa(\tau)-t^{2}\kappa(\tau))(1-t^{2})(1+tg\widetilde{Y})}\right)\\ &+\frac{\widetilde{X}-\widetilde{Y}}{2t}\left(\mu^{2}(\tau)\kappa^{\prime}(\tau)\tau_{z}+\frac{f_{z}}{f(1+tg\widetilde{Y})}\right)+\frac{f\sqrt{1-t^{2}}}{1+tg\widetilde{Y}}\left(-\frac{2t^{2}\sqrt{1-t^{2}}\widetilde{X}\widetilde{Y}}{\tau\kappa(\tau)c}+\widetilde{X}t[\kappa^{\prime}(\tau)(2t^{2}-1)+\tau\kappa^{\prime\prime}(\tau)]\right)\tau_{z}\\ &+\frac{f_{z}\sqrt{1-t^{2}}}{(1+tg\widetilde{Y})^{2}}\left(-\frac{2t^{2}\sqrt{1-t^{2}}\widetilde{X}\widetilde{Y}}{c}+\widetilde{X}t[\tau\kappa^{\prime}(\tau)-1-2\kappa(\tau)+t^{2}2\kappa(\tau))]\right)\\ &+\frac{fg\widetilde{Y}\sqrt{1-t^{2}}}{c(1+tg\widetilde{Y})^{2}}\left(\frac{\widetilde{X}-\widetilde{Y}}{2\mu^{2}(\tau)}-\frac{2t^{3}\sqrt{1-t^{2}}\widetilde{X}\widetilde{Y}}{c}+\widetilde{X}t^{2}[\tau\kappa^{\prime}(\tau)-1-2\kappa(\tau)+2t^{2}\kappa(\tau))]\right)c_{z},\\ g_{1}(z,t)=&\frac{1}{2t}+\frac{(\kappa(\tau)+1)g\widetilde{X}+t(1+2\kappa(\tau)-t^{2}\kappa(\tau))(1-tg\widetilde{X})}{2(1-t^{2})(1-tg\widetilde{X})(1+\kappa(\tau)-t^{2}\kappa(\tau))}\\ &+\frac{f\sqrt{1-t^{2}}}{1-tg\widetilde{X}}\left(t(\tau\kappa^{\prime}(\tau)-1-2\kappa(\tau)+2t^{2}\kappa(\tau))+\frac{2t^{2}\sqrt{1-t^{2}}\widetilde{X}}{c}\right),\\ g_{2}(z,t)=&-\frac{1}{2t}-\frac{(\kappa(\tau)+1)g\widetilde{X}+t(1+4\kappa(\tau)-2t^{2}\kappa(\tau))(1-tg\widetilde{X})}{2(1-t^{2})(1-tg\widetilde{X})(1+\kappa(\tau)-t^{2}\kappa(\tau))}+\frac{2t^{2}f(1-t^{2})\widetilde{Y}}{c(1-tg\widetilde{X})}\\ &-\frac{fg\sqrt{1-t^{2}}}{(1-tg\widetilde{X})^{2}}\left(\frac{\widetilde{X}-\widetilde{Y}}{2\mu^{2}(\tau)}-(\tau\kappa^{\prime}(\tau)-1-2\kappa(\tau)+2t^{2}\kappa(\tau))t^{2}\widetilde{Y}-\frac{2t^{3}f\sqrt{1-t^{2}}\widetilde{X}\widetilde{Y}}{c}\right),\\ g_{3}(z,t)=&(\widetilde{Y}-\widetilde{X})\left(\kappa^{\prime}(\tau)\mu^{2}(\tau)\tau_{z}+\frac{f_{z}}{f(1-tg\widetilde{X})}\right)\left(\frac{(\kappa(\tau)+1)g\widetilde{X}+t(1+4\kappa(\tau)-2t^{2}\kappa(\tau))(1-tg\widetilde{X})}{2(1+\kappa(\tau)-t^{2}\kappa(\tau))(1-t^{2})(1-tg\widetilde{X})}\right)\\ &+\frac{\widetilde{Y}-\widetilde{X}}{2t}\left(\kappa^{\prime}(\tau)\mu^{2}(\tau)\tau_{z}+\frac{f_{z}}{f(1-tg\widetilde{X})}\right)+\frac{f\sqrt{1-t^{2}}}{1-tg\widetilde{X}}\left(\frac{2t^{2}\sqrt{1-t^{2}}\widetilde{X}\widetilde{Y}}{\tau\kappa(\tau)c}+\widetilde{Y}t[\kappa^{\prime}(\tau)(2t^{2}-1)+\tau\kappa^{\prime\prime}(\tau)]\right)\tau_{z}\\ &+\frac{f_{z}\sqrt{1-t^{2}}}{(1-tg\widetilde{X})^{2}}\left(\frac{2t^{2}\sqrt{1-t^{2}}\widetilde{X}\widetilde{Y}}{c}+\widetilde{Y}t[\tau\kappa^{\prime}(\tau)-1-2\kappa(\tau)+t^{2}2\kappa(\tau))]\right)\\ &+\frac{fg\widetilde{X}\sqrt{1-t^{2}}}{c(1-tg\widetilde{X})^{2}}\left(\frac{\widetilde{X}-\widetilde{Y}}{2\mu^{2}(\tau)}-\frac{2t^{3}\sqrt{1-t^{2}}\widetilde{X}\widetilde{Y}}{c}-\widetilde{Y}t^{2}[\tau\kappa^{\prime}(\tau)-1-2\kappa(\tau)+2t^{2}\kappa(\tau))]\right)c_{z}.\end{split}$
@@
$\begin{split}\tilde{\partial}_{+}\Xi=\left(\frac{5}{2t}+\hat{f}_{1}(z,t)\right)\Xi+\hat{f}_{2}(z,t)\frac{\Theta}{2t}+\hat{f}_{3}(z,t)\frac{t}{2},\\ \tilde{\partial}_{-}\Theta=\left(\frac{5}{2t}+\hat{g}_{1}(z,t)\right)\Theta+\hat{g}_{2}(z,t)\frac{\Xi}{2t}+\hat{g}_{3}(z,t)\frac{t}{2},\end{split}$ (3.36)
@@
$\begin{split}\hat{f}_{1}&=h+f_{1}-\frac{1}{2t},\;\hat{f}_{2}=2tf_{2},\;\hat{f}_{3}=\frac{2}{t}(\tilde{\Lambda}_{+}-\tilde{\Lambda}_{-})f_{3},\\ \hat{g}_{1}&=h+g_{1}-\frac{1}{2t},\;\hat{g}_{2}=2tg_{2},\;\hat{g}_{3}=\frac{2}{t}(\tilde{\Lambda}_{+}-\tilde{\Lambda}_{-})g_{3}.\end{split}$
@@
$\begin{split}\left(\frac{\hat{c}\hat{f}t}{t\hat{g}-\hat{b}}-\frac{\hat{c}\sqrt{1+(\psi^{\prime})^{2}}}{\hat{\bar{\omega}}^{2}\hat{\bar{\omega}}^{\prime}}\right)\cdot\tilde{X}_{z}|_{\widehat{M^{\prime}P^{\prime}}}=&\frac{\hat{b}+\hat{a}}{2t^{2}\hat{a}\hat{b}}+\frac{\hat{b}+\hat{a}}{\hat{a}\hat{b}t}\hat{H}_{11}+\hat{H}_{12}+\frac{\hat{d}}{t\hat{a}^{2}}-\frac{t\hat{d}^{\prime}}{\hat{\bar{\omega}}\hat{\bar{\omega}}^{\prime}\hat{a}^{2}}-\frac{1}{\hat{\bar{\omega}}\hat{\bar{\omega}}^{\prime}\hat{a}^{2}}\left(\frac{\cos\hat{\theta}\psi^{\prime\prime}}{\kappa(\hat{\tau})\bar{\omega}(1+(\psi)^{\prime 2})}\right)^{\prime}\\ =&\frac{2\hat{d}^{2}}{\hat{a}^{2}\hat{b}}+\frac{2\hat{d}}{\hat{a}\hat{b}}\hat{H}_{11}+\hat{H}_{12}-\frac{t\hat{d}^{\prime}}{\hat{\bar{\omega}}\hat{\bar{\omega}}^{\prime}\hat{a}^{2}}-\frac{1}{\hat{\bar{\omega}}\hat{\bar{\omega}}^{\prime}\hat{a}^{2}}\left(\frac{\cos\hat{\theta}\psi^{\prime\prime}}{\kappa(\hat{\tau})\bar{\omega}(1+(\psi)^{\prime 2})}\right)^{\prime},\end{split}$
@@
$\begin{split}\tilde{\partial}_{+}\left(t^{-\frac{1}{2}-\nu}\widetilde{\Xi}\right)&=t^{-\frac{3}{2}-\nu}\left(\hat{f}_{1}t\widetilde{\Xi}+\frac{\widetilde{\Theta}}{2}\hat{f}_{2}+\frac{t^{\nu}}{2}\hat{f}_{3}\right),\\ \tilde{\partial}_{-}\left(t^{-\frac{1}{2}-\nu}\widetilde{\Theta}\right)&=t^{-\frac{3}{2}-\nu}\left(\hat{g}_{1}t\widetilde{\Theta}+\frac{\widetilde{\Xi}}{2}\hat{g}_{2}+\frac{t^{\nu}}{2}\hat{g}_{3}\right).\end{split}$ (3.39)